# Chemical Diagnostics to Unveil Environments Enriched by First Stars

### v1

@@ Abstract
Unveiling the chemical fingerprints of the first (Pop III) stars is crucial for indirectly studying their properties and probing their massive nature. In particular, very massive Pop III stars explode as energetic Pair-Instability Supernovae (PISNe), so their chemical products might escape in the diffuse medium around galaxies, opening the possibility to observe their fingerprints in distant gas clouds. Recently, three $z>6.3$ absorbers with abundances consistent with an enrichment from PISNe have been observed with JWST. In this Letter, we present novel chemical diagnostics to uncover environments mainly imprinted by PISNe. Furthermore, we revise the JWST low-resolution measurements by analysing the publicly available high-resolution X-Shooter spectra for two of these systems. Our results reconcile the chemical abundances of these absorbers with those from literature, which are found to be consistent with an enrichment dominated ($>50\%$ metals) by normal Pop II SNe. We show the power of our novel diagnostics in isolating environments uniquely enriched by PISNe from those mainly polluted by other Pop III and Pop II SNe. When the subsequent enrichment from Pop II SNe is included, however, we find that the abundances of PISN-dominated environments partially overlap with those predominantly enriched by other Pop III and Pop II SNe. We dub these areas confusion regions. Yet, the odd-even abundance ratios [Mg,Si/Al] are extremely effective in pinpointing PISNe-dominated environments and allowed us to uncover, for the first time, an absorber consistent with a PISN enrichment for all the six measured elements.
@@ 1 Introduction
Pop III SNe with various progenitor masses and explosion energies synthesize different heavy elements (e.g. Heger & Woosley, 2010; Kobayashi, 2012) thus producing distinctive chemical fingerprints (see Vanni et al., 2023a). These unique signatures can be observed in distant gaseous absorbers directly enriched by these pristine sources (Saccardi et al., 2023), or in long-lived metal-poor stars born in Pop III enriched environments (e.g. Frebel et al., 2019; Skúladóttir et al., 2021). Stellar evolution calculations for the chemical products of very massive pair instability SNe (PISNe, $\rm M_{*}=140-260M_{\odot}$) are extremely robust and predict unique abundance patterns, featuring a strong odd-even effect (Heger & Woosley, 2002; Takahashi, 2018).
@@
PISNe are very energetic ($\rm E_{SN}\in[10^{52};10^{53}]$ erg), thus their metals can easily escape their hosting galaxies (e.g. Bromm et al., 2001; Smith et al., 2015) filling the diffuse circum-galactic and inter-galactic media (e.g. Pallottini et al., 2014). Thus, the long-waited discovery of this PISN descendant opens the concrete possibility to pinpoint the chemical fingerprint of PISNe in high-$z$ diffuse absorbers. Very recently, Christensen et al. (2023) reported chemical abundances consistent with a PISN enrichment for three $z>6.3$ absorbers observed in JWST QSO’s spectra. Indeed, they found that these systems have high [Si/O]$>0.5$ values and only upper limits for [C/O]$<0$ (see e.g. Ma et al., 2017; D’Odorico et al., 2023a).
@@ 2 Summary of the model
For this study we use the model first introduced by Salvadori et al. (2019) and then generalised by Vanni et al. (2023b). This simple and general parametric study investigates the chemical abundances (elements from C to Zn) of gaseous environments imprinted by a single Pop III SN and subsequently polluted by normal (Pop II) SNe. Pop III stars are indeed expected to predominantly form in isolation in low-mass mini-halos (e.g. Hirano et al., 2015). Furthermore, simulations show that in these proto-galaxies self-enrichment from a single Pop III SN can promptly trigger the formation of normal Pop II stars by enriching the inter-stellar medium (ISM) above the critical metallicity value (e.g. Ritter et al., 2012). The model is quite robust since our findings for ancient metal-poor stars (Vanni et al., 2023b; Skúladóttir et al., 2024) have been confirmed by more sophisticated cosmological chemical evolution models and simulations (Rossi et al., 2023; Koutsouridou et al., 2023). The model was used to unveil the Pop III star’s signature in distant gaseous absorbers (Salvadori et al., 2023, Sodini et al. submitted), which is a proof of its versatility.
@@
The innovative aspects of our approach with respect to other simple models available in literature (e.g. Ishigaki et al., 2018; Welsh et al., 2021) are reported below.
@@
A parametric study. The uncertainties related to early cosmic star-formation are enclosed in two free parameters: the star-formation efficiency, $f_{*}$, which quantifies the amount of cold gas turned into stars, and the dilution factor, $f_{\rm dil}$, which varies depending upon the amount of gas available to dilute metals. A third parameter, $f_{\rm PopIII}$, sets the fraction of metals in the ISM provided by Pop III with respect to Pop II SNe. The chemical abundances of the gas, [X/H], depend on the ratio $f_{*}/f_{\rm dil}$ and $f_{\rm PopIII}$ (see Vanni et al., 2023b), and thus are evaluated by varying these unknowns in the full parameter space $f_{*}/f_{\rm dil}\in[10^{-4};10^{-1}]$ and $f_{\rm PopIII}\in[1,0.01]$. Conversely, the relative abundance ratios of different chemical elements, [X/Y], are strongly influenced by $f_{\rm PopIII}$ but they only indirectly depend on $f_{*}/f_{\rm dil}$, which sets the initial metallicity of subsequent generations of Pop II stars (see Salvadori et al., 2019).
@@
Exploring the properties of Pop III SNe. To investigate the whole range of unknown properties of Pop III SNe, we exploit the tabulated yields of Heger & Woosley (2010), which represent the most complete datasets for Pop III SNe with progenitor masses between $\rm 10-100\leavevmode\nobreak\ M_{\odot}$. Indeed, the authors investigate the chemical products of these SNe, that can potentially explode with different energies: from the lowest-energy faint SNe, $\rm E_{SN}=0.6\times 10^{51}$ erg, to the most energetic hypernovae, $\rm E_{SN}=10^{52}$ erg, adopting different mixing parameters. Here we explore this full parameter space. For more massive, $\rm m_{\star}=140-260M_{\odot}$, Pop III stars exploding as energetic PISNe, $\rm E_{SN}=10^{52}-10^{53}$ erg, we adopt the yields by Heger & Woosley (2002).
@@
Subsequent pollution by Pop II SNe. To study how the chemical abundances vary due to the further contribution of Pop II SNe, we assume that Pop II stars are distributed according to a Larson’s IMF, $\phi(m_{\star})\propto m_{\star}^{-2.35}\times\exp{\left(-{m_{\rm ch}}/{m_{\star}}\right)}$ with $m_{\rm ch}=0.35\rm M{\odot}$ and compute the contribution of Pop II stars above different masses, which allow us to evaluate the effect of SNe evolving on different time-scales (see Salvadori et al., 2019). For Pop II SNe we adopt the recent yields from Limongi & Chieffi (2018, non-rotating, set R).
@@ 3 Observational data
The two high-redshift QSOs UHS J0439+1634 ($z_{\rm em}=6.5185$) and ULAS J1342+0928 ($z_{\rm em}=7.5353$) have been observed recently with JWST/NIRSpec at low spectral resolution ($\rm R\approx 2700$) and their spectra were analysed by Christensen et al. (2023). Both objects also have intermediate resolution ($\rm R\approx 9000$) good signal-to-noise ratio VLT/X-Shooter data. The X-Shooter spectrum of UHS J0439+1634 was presented in D’Odorico et al. (2023b) and analysed in Davies et al. (2023). However, these authors do not identify the $\scriptstyle\rm I$ 1302 Å̊$\mathring{\rm A}$ transition associated with the system at $z=6.2897$, discussed by Christensen et al. (2023), because it falls in a spectral region already populated by other absorption lines (see Davies et al., 2023, for details). Our analysis of the system is reported in Appendix A.
@@
The analysis of the X-shooter absorption spectrum of J1342+0928 is carried out in this work for the first time, and is briefly described in Appendix A. For the purpose of this paper, we analysed only the two $\scriptstyle\rm I$ absorption systems at $z=7.369$ and $z=7.443$ for which Christensen et al. (2023) reported the chemical abundances. However, we carried out a first look analysis of the whole spectrum to verify that the transitions of the considered systems were not blended with lines from other systems.
@@
Observed absorption features were identified and fitted with Voigt profiles to determine ionic column densities. Assuming that these systems are metal-poor DLAs, we derived chemical abundances from the ratio of the measured column densities (see e.g. Becker et al., 2012, Sodini et al. subm.) and adopting solar chemical abundances from Asplund et al. (2009). Measured column densities and plots of the absorption lines are reported in Appendix A.
3.1 New results
@@
The abundances for the system in UHS J0439+1634 and the 2 systems in ULAS J1342+0928 as derived by Christensen et al. (2023) are shown in Fig. 1 with light red points. These points are connected with dashed lines to the revised abundances derived in this work (red stars), which are reported in Table 1. The new abundances derived from X-Shooter data agree with other systems in the literature, thanks to a decrease of the [Si/O] values of these absorbers of $\gtrsim 0.5$ dex. The changes in relative abundances are generally due to an overestimate of the Si $\scriptstyle\rm II$ column densities in the NIRSpec spectra, possibly deriving from line blending. The only point from Christensen et al. (2023) that remains consistent with an enrichment from PISNe is due to the system at $z=6.5625$ toward VDES J0020-3653, which at present does not have sufficient high-resolution observations to carry out the analysis of absorption lines.
@@
Figure 1: Chemical abundances predicted for environments uniquely enriched by PISNe (red), by other Pop III SNe (faint SNe: green; core-collapse SNe: blue; hypernovae: purple), and by normal Pop II SNe (grey shaded areas). Each environment is enriched by a single Pop III SN with different mass and explosion energy. The contours with decreasing color intensity identify 30%, 60% and 90% probability densities, corresponding to the fraction of environments predicting those abundances for each SN type. Points with errorbars are the measured abundances of absorbers in the literature (Sec. 1, labels). The values reported by Christensen et al. (2023, red) are connected with the new measurements obtained using higher-resolution spectra (star). The absorber in the sightline to J1018+0548 (Saccardi et al., 2023, $z=3.39$) is highlighted with a black cross (see Sec. 4.2). The arrows show in which direction dust depletion would modify the predicted abundances.
@@ 4.1 Purely Pop III enriched environments
In Fig. 1 (top right), we see that only one of the systems presented in Christensen et al. (2023) remains consistent with an imprint from PISN. If the [Si/O] JWST measurement will be confirmed with higher resolution data, we predict the following ranges of abundances for other elements, which would imply that this absorber was truly and uniquely polluted by PISNe: [O/Fe]$\simeq[-0.35;1.04]$, [Al/Si]$\simeq[-2.16;-1.95]$, [Si/Fe]$\simeq[0.40;1.58]$ and [Al/Mg]$\simeq[-1.38;-1.28]$.
@@ 4.2 Pop II contribution: the confusion region
In Fig. 2 we show how the injection of heavy elements from Pop II SNe changes the chemical abundances of PISN- and other Pop III SN pre-enriched environments. To this end, we compute the probability density regions of environments solely ($100\%$ metals), mainly ($\geq 60\%$), and only partially ($\leq 50\%$) enriched by Pop III (PISNe or other SNe), thus varying $f_{\rm PopIII}$. We see that the contribution of normal Pop II SNe enlarges the regions predominantly imprinted by PISNe, which partially overlap with those mainly enriched by other Pop III/II SNe. Furthermore, while there is still a partial separation between environments mainly polluted by PISNe and by other Pop III SNe, when Pop II SNe become the dominant contributors the two regions almost completely overlap, meaning that we are loosing the information about the pristine SNe that initially enriched these media.
In each diagnostic of Fig. 2, we notice the emergence of regions in which the chemical abundance ratios predicted for Pop III- (PISNe and other pristine SNe) and Pop II-dominated environments blend. We dub this area the confusion region, since environments dwelling in this specific zone might have been imprinted by very different sources: from a single Pop III SN, to a population of normal Pop II SNe. In Fig. 2, we see that the majority of the absorption systems present in the literature reside within the confusion regions and in particular, in most of the proposed diagnostics, they are consistent with an enrichment dominated by normal Pop II SNe, but not solely driven by Pop II SNe (see Fig. 1).
To pinpoint PISN- and other Pop III-enriched environments, we should thus look for absorbers that dwell outside of the confusion regions, i.e. in areas where the Pop III chemical fingerprints are preserved. To this end, we see that the odd-even abundance ratios are extremely effective in locating PISN dominated environments, for example [Al/Mg] $<-0.75$ and [Al/Si] $<-1.5$. The absorption system at $z=3.39$ in the spectrum of J1018+0548 from Saccardi et al. (2023, black cross in Figs. 1,2,3) lies at [Al/Mg] $=-1.23$ and [Si/Fe] $=0.13$ implying that it might be predominantly enriched by PISNe. However, when we consider its positions in other diagnostic plots, we see that this absorber is also consistent with being predominantly enriched by both PISNe and other Pop III SNe.

### v2

@@ Abstract
Unveiling the chemical fingerprints of the first (Pop III) stars is crucial for indirectly studying their properties and probing their massive nature. In particular, very massive Pop III stars explode as energetic Pair-Instability Supernovae (PISNe), allowing their chemical products to escape in the diffuse medium around galaxies, opening the possibility to observe their fingerprints in distant gas clouds. Recently, three $z>6.3$ absorbers with abundances consistent with an enrichment from PISNe have been observed with JWST. In this Letter, we present novel chemical diagnostics to uncover environments mainly imprinted by PISNe. Furthermore, we revise the JWST low-resolution measurements by analysing the publicly available high-resolution X-Shooter spectra for two of these systems. Our results reconcile the chemical abundances of these absorbers with those from literature, which are found to be consistent with an enrichment dominated ($>50\%$ metals) by normal Pop II SNe. We show the power of our novel diagnostics in isolating environments uniquely enriched by PISNe from those mainly polluted by other Pop III and Pop II SNe. When the subsequent enrichment from Pop II SNe is included, however, we find that the abundances of PISN-dominated environments partially overlap with those predominantly enriched by other Pop III and Pop II SNe. We dub these areas confusion regions. Yet, the odd-even abundance ratios [Mg,Si/Al] are extremely effective in pinpointing PISN-dominated environments and allowed us to uncover, for the first time, an absorber consistent with a combined enrichment by a PISN and another Pop III SN for all the six measured elements.
@@ 1 Introduction
Pop III SNe with various progenitor masses and explosion energies synthesize different heavy elements (e.g. Heger & Woosley, 2010; Kobayashi, 2012) thus producing distinctive chemical fingerprints (see Vanni et al., 2023a). These unique signatures can be observed in distant gaseous absorbers directly enriched by these pristine sources (Saccardi et al., 2023), or in long-lived metal-poor stars born in Pop III enriched environments (e.g. Frebel et al., 2019; Skúladóttir et al., 2021). Stellar evolution calculations for the chemical products of very massive pair instability SNe (PISNe, $\rm M_{*}=140-260M_{\odot}$, Barkat et al., 1967) are extremely robust and predict unique abundance patterns, featuring a strong odd-even effect (Heger & Woosley, 2002; Takahashi, 2018).
@@
PISNe are very energetic ($\rm E_{SN}\in[10^{52};10^{53}]$ erg), thus their metals can easily escape their hosting galaxies (e.g. Bromm et al., 2001; Smith et al., 2015) filling the diffuse circum-galactic and inter-galactic media (e.g. Pallottini et al., 2014). Thus, the long-waited discovery of this PISN descendant opens the concrete possibility to pinpoint the chemical fingerprint of PISNe in high-$z$ diffuse absorbers too. Very recently, Christensen et al. (2023) reported chemical abundances consistent with a PISN enrichment for three $z>6.3$ absorbers observed in JWST QSO’s spectra. Indeed, they found that these systems have high [Si/O]$>0.5$ values and only upper limits for [C/O]$<0$ (see e.g. Ma et al., 2017a; D’Odorico et al., 2023a).
@@ 2 Summary of the model
For this study, we use the model first introduced by Salvadori et al. (2019) and then generalised by Vanni et al. (2023b). This simple and general parametric study investigates the chemical abundances (elements from C to Zn) of gaseous environments imprinted by a single Pop III SN and subsequently polluted by normal (Pop II) SNe. Pop III stars are indeed expected to predominantly form in isolation in low-mass mini-halos (e.g. Hirano et al., 2015). Furthermore, simulations show that in these proto-galaxies self-enrichment from a single Pop III SN can promptly trigger the formation of normal Pop II stars by enriching the inter-stellar medium (ISM) above the critical metallicity value (e.g. Ritter et al., 2012). The model is quite robust since our findings for ancient metal-poor stars (Vanni et al., 2023b; Skúladóttir et al., 2024) have been confirmed by more sophisticated cosmological chemical evolution models and simulations (Rossi et al., 2023; Koutsouridou et al., 2023). The model was used to unveil the Pop III star’s signature in distant gaseous absorbers (Salvadori et al., 2023, Sodini et al. submitted), which is a proof of its versatility.
@@
The innovative aspects of our approach with respect to other simple models available in literature (e.g. Ishigaki et al., 2018; Welsh et al., 2019) are reported below.
@@
A parametric study. The uncertainties related to early cosmic star-formation are enclosed in two free parameters: the star-formation efficiency, $f_{*}$, which quantifies the amount of cold gas turned into stars, and the dilution factor, $f_{\rm dil}$, which varies depending upon the amount of gas available to dilute metals. A third parameter, $f_{\rm PopIII}$, sets the fraction of metals in the ISM provided by Pop III with respect to Pop II SNe. The chemical abundances of the gas, [X/H], depend on the ratio $f_{*}/f_{\rm dil}$ and $f_{\rm PopIII}$ (see Vanni et al., 2023b), and thus are evaluated by varying these unknowns in the full parameter space: $f_{*}/f_{\rm dil}\in[10^{-4};10^{-1}]$ and $f_{\rm PopIII}\in[1,0.01]$. Conversely, the relative abundance ratios of different chemical elements, [X/Y], are strongly influenced by $f_{\rm PopIII}$ but they only indirectly depend on $f_{*}/f_{\rm dil}$, which sets the initial metallicity of subsequent generations of Pop II stars (see Salvadori et al., 2019).
@@
Exploring the properties of Pop III SNe. To investigate the whole range of unknown properties of Pop III SNe, we exploit the tabulated yields of Heger & Woosley (2010), which represent the most complete datasets for Pop III SNe with progenitor masses between $\rm 10-100\leavevmode\nobreak\ M_{\odot}$. Indeed, the authors investigate the chemical products of these SNe, that can potentially explode with different energies: from the lowest-energy faint SNe, $\rm E_{SN}=0.6\times 10^{51}$ erg, to the most energetic hypernovae, $\rm E_{SN}=10^{52}$ erg, adopting different mixing parameters. Here we explore this full parameter space. For more massive, $\rm m_{\star}=140-260M_{\odot}$, Pop III stars exploding as energetic PISNe, $\rm E_{SN}=10^{52}-10^{53}$ erg, we adopt the yields by Heger & Woosley (2002). For our exploration we are thus implicitly assuming that Pop III SNe with different progenitor mass within the range $\rm m_{\star}=10-100M_{\odot}$ and $\rm m_{\star}=140-260M_{\odot}$, and different SN explosion energies have an equal probability to explode. Note that considering a burst of Pop III SNe with the same mass/explosion energy instead of a single SN would not change the results. Finally, here we also investigate how the abundance ratios of PISNe-enriched environments change after the explosion of a second Pop III SN (see Appendix B).
@@
Subsequent pollution by Pop II SNe. To study how the chemical abundances vary due to the further contribution of Pop II SNe, we assume that Pop II stars are distributed according to a Larson’s IMF, $\phi(m_{\star})\propto m_{\star}^{-2.35}\times\exp{\left(-{m_{\rm ch}}/{m_{\star}}\right)}$ with $m_{\rm ch}=0.35\rm M_{\odot}$ and compute the contribution of Pop II stars above different masses, which allow us to evaluate the effect of SNe evolving on different time-scales (see Salvadori et al., 2019). For Pop II SNe, we adopt the recent yields from Limongi & Chieffi (2018, non-rotating, set R).
@@ 3 Observational data
3.1 New measurements
The analysis of the three systems with possible PISN enrichment by Christensen et al. (2023) have been carried out in JWST/NIRSpec spectra at low spectral resolution ($\rm R\approx 2700$). Two of these systems fall in the spectra of the QSOs UHS J0439+1634 ($z_{\rm em}=6.5185$) and ULAS J1342+0928 ($z_{\rm em}=7.5353$), respectively, which also have intermediate resolution ($\rm R\approx 9000$) good signal-to-noise ratio VLT/X-Shooter data.
@@ 3.1 New measurements
The X-Shooter spectrum of J0439+1634 was presented in D’Odorico et al. (2023b) and analysed in Davies et al. (2023). However, these authors do not identify the $\scriptstyle\rm I$ 1302 Å̊$\mathring{\rm A}$ transition associated with the system at $z=6.2897$, discussed by Christensen et al. (2023), because it falls in a spectral region already populated by other absorption lines (see Davies et al., 2023, for details). Our analysis of the system is reported in Appendix A.
The analysis of the X-shooter absorption spectrum of J1342+0928 is carried out in this work for the first time, and is described in Appendix A.
@@
For the purpose of this paper, we analysed only the two $\scriptstyle\rm I$ absorption systems at $z=7.369$ and $z=7.443$ for which Christensen et al. (2023) reported the chemical abundances, the former showing possible PISN enrichment. However, we carried out a first look analysis of the whole spectrum to verify that the transitions of the considered systems were not blended with lines from other systems.
@@
Observed absorption features were identified and fitted with Voigt profiles to determine ionic column densities. We derived chemical abundances from the ratio of the measured column densities and adopting solar chemical abundances from Asplund et al. (2009). Based on the common hypothesis that these systems are metal-poor DLAs (e.g. Becker et al., 2012; Christensen et al., 2023), we did not correct for ionization effects and dust depletion. Measured column densities and plots of the absorption lines are reported in Appendix A.
@@
The abundances for the system in UHS J0439+1634 and the 2 systems in ULAS J1342+0928 as derived by Christensen et al. (2023) are shown in Fig. 1 with red empty points. These points are connected with dashed black lines to the revised abundances derived in this work (red stars), which are reported in Table 1. The new abundances derived from X-Shooter data agree with other systems in the literature, thanks to a decrease of the [Si/O] values of these absorbers of $\gtrsim 0.4$ dex. The changes in relative abundances are generally due to an overestimate of the Si $\scriptstyle\rm II$ column densities in the NIRSpec spectra, possibly deriving from line blending. The only point from Christensen et al. (2023) that remains consistent with an enrichment from PISNe is due to the system at $z=6.5625$ toward VDES J0020-3653, which at present does not have sufficient high-resolution observations to carry out the analysis of absorption lines.
@@
Figure 1: Chemical abundances predicted for environments uniquely enriched by PISNe (red), by other Pop III SNe (faint SNe: green; core-collapse SNe: blue; hypernovae: purple), and by normal Pop II SNe (grey shaded areas). Each environment is enriched by a single Pop III SN with different mass and explosion energy. The contours with decreasing color intensity identify 30%, 60% and 90% probability densities, corresponding to the fraction of environments predicting those abundances for each SN type. Points with errorbars are the measured abundances of absorbers in the literature (Sec. 1, labels). The values reported by Christensen et al. (2023, red empty) are connected with the new measurements obtained using higher-resolution spectra (star). The absorber in the sightline to J1018+0548 (Saccardi et al., 2023, $z=3.39$) is highlighted with a black cross (see Sec. 4.2). The arrows show in which direction dust depletion would modify the predicted abundances.
@@ 4.1 Purely Pop III enriched environments
In Fig. 1 (top right), we see that only the system toward VDES J0020-3653 in Christensen et al. (2023) remains consistent with an imprint from PISN. Comparing the [Si/O] value of this absorber with the predictions of our model for the environments polluted solely by PISN, we determine the possible initial masses for the PISN that enriched this absorber, $m_{*}\in[183.0;234.0]\rm M_{\odot}$. Then, we predict the range of abundances for other elements, which would imply that this absorber was truly and uniquely polluted by PISNe: [O/Fe]$\simeq[-0.35;1.04]$, [Al/Si]$\simeq[-2.16;-1.95]$, [Si/Fe]$\simeq[0.40;1.58]$ and [Al/Mg]$\simeq[-1.38;-1.28]$.
@@ 4.2 Pop II contribution: the confusion region
In Fig. 2 we show how the injection of heavy elements from Pop II SNe changes the chemical abundances of PISN and other Pop III SN pre-enriched environments. To this end, we compute the probability density regions of environments solely ($100\%$ metals), mainly ($\geq 60\%$), and only partially ($\leq 50\%$) enriched by Pop III (PISNe or other SNe), thus varying $f_{\rm PopIII}$. We see that the contribution of normal Pop II SNe enlarges the regions predominantly imprinted by PISNe, which partially overlap with those mainly enriched by other Pop III/II SNe. Furthermore, while there is still a partial separation between environments mainly polluted by PISNe and by other Pop III SNe, when Pop II SNe become the dominant contributors the two regions almost completely overlap, meaning that we are loosing the information about the pristine SNe that initially enriched these media.
In each diagnostic of Fig. 2, we notice the emergence of regions in which the chemical abundance ratios predicted for Pop III- (PISNe and other pristine SNe) and Pop II-dominated environments blend. We dub this area the confusion region, since environments dwelling in this specific zone might have been imprinted by different sources: one Pop III SN (or two, see Appendix B), a mixture of Pop III and Pop II SNe or a population of normal Pop II SNe. In Fig. 2, we see that the majority of the absorption systems present in the literature reside within the confusion regions and in particular, in most of the proposed diagnostics, they are consistent with an enrichment dominated by normal Pop II SNe, but not solely driven by Pop II SNe (see Fig. 1). Thus, the nature of the absorbers dwelling in the confusion regions might remain undetermined, even by studying their abundance patterns with the currently available $\leq 5$ abundance ratios.
To pinpoint PISN- and other Pop III-enriched environments, we should thus look for absorbers that dwell outside of the confusion regions, i.e. in areas where the Pop III chemical fingerprints are preserved. To this end, we see that the odd-even abundance ratios are extremely effective in locating PISN dominated environments, for example [Al/Mg] $<-0.75$ and [Al/Si] $<-1.5$ (see also Takahashi et al., 2018). The absorption system at $z=3.39$ in the spectrum of J1018+0548 from Saccardi et al. (2023, black cross in Figs. 1,2,3) lies at [Al/Mg] $=-1.23$ and [Si/Fe] $=0.13$ implying that it might be predominantly enriched by PISNe. However, when we consider its positions in other diagnostic plots, we see that this absorber is also consistent with being predominantly enriched by both PISNe and other Pop III SNe.